OpenStudy (gabylovesyou):

@hartnn Megan and Julie are stuck simplifying radical expressions. Megan has to simplify the quantity of x to the one third power, over x to the one twelfth power. Julie has to simplify the thirty second root of the quantity of x times x to the second times x to the fifth. Using full sentences, describe how to fully simplify Megan and Julie's expressions. Describe if Megan and Julie started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (mathmale):

If you're dividing two powers of x, as in \[\frac{ x^a }{ x^b }\]

OpenStudy (mathmale):

... the result is \[\frac{ x^a }{ x^b }=x ^{a-b}\]

OpenStudy (mathmale):

Use this rule to evaluate Megan's quotient.

OpenStudy (mathmale):

Yes, but that's not the complete answer you want. First of all, y ou need to keep that base, x, in your expression. Secondly, you must find the LCD and then combine (1/3) and (1/12).

OpenStudy (mathmale):

sorry, but no, it's not 3.

OpenStudy (mathmale):

I'm sure you've added fractions before. Come on!\[\frac{ 1 }{ 3 }-\frac{ 1 }{ 12}\]

OpenStudy (mathmale):

By what integer should you multiply the numerator and denominator of 1/3 to obtain '12' in the denominator?

OpenStudy (mathmale):

Do that multiplication now, please.

OpenStudy (mathmale):

\[\frac{ x^a }{ x^b }=x ^{a-b}=x ^{\frac{ ? }{ 12 }}\]

OpenStudy (mathmale):